OpenStudy (anonymous):

cot(theta)=3/4 and the terminal point determined by theta is in quadrant 3, then: Please select all correct answers and click submit. a. tan(theta)=4/3 b.csc(theta)=-5/3 c.sintheta=3/5 d.costheta=-3/5

OpenStudy (jennychan12):



OpenStudy (jennychan12):

that would mean that tan θ = 4/3. sin θ = -4/5. cos θ = -3/5, csc θ = -5/4, sec θ = -5/3 that would be A and D

OpenStudy (jennychan12):

@kaykim95 do u get it? o.O